how many gallons of 25% sugar content kool aid and 40% sugar content kool aid need to be mixed in order to make 50 gallons of 30% sugar content kool aid?
Answer by chelcilovesmj1(3)

Let
A= the amount of 25% kool aid
B= the amount of 40% kool aid
2 equations are needed
1.) A+B = 50 gallons of kool aid
2.) .25(A) + .40(B) = .30(50)
1.)rewrite 1st eqaution as such... B = 50-A
2.)substitute newly written equation as such...
.25(A) + .40(50-A) = 15
.25(A) - .40(A) + 20 = 15
-.15A = -5
A = (approx) 33.3 gallons
3.) use 1st equation to solve for B
33.3 + B = 50
B = 16.7 gallons
4.) Final Answer
A = 33.3 gallons, B = 16.7 gallons
hope this helps :)